Identify the variation as direct, inverse, joint or combined.
m/n=4

Answer by Theo(13342)

m/n = 4 looks like it is direct.
multiply both sides by n and you get:
m = 4n
the 4 could represent k in the formula m = kn which is a direct variation formula.
if it was direct, then an increase in m would correleate to an increase in n.
let's see if that's true.
allow m = 1
you get 1/n = 4 and you solve for n to get n = 1/4
your ratio becomes 1/(1/4) = 4 which is true.
now allow m = 2
you get 2/n = 4 and you solve for n to get n = 2/4
your ratio becomes 2/(2/4) = 4 which is also true.
as m increased, n increases, so the variation is direct.
allow m to be 100.
the formua says that 100/n = 4
multiply both sides by n to get 100 = 4n
divide both sides by 4 to get n = 25
your ratio is 100/25 = 4 which is true
as m increased, n increased.
direct variation again.
direct is m = kn which leads to m/n = k
inverse is m = k/n which leads to mn = k
joint is m = kn/p which leads to k = mp/n
i'd go for direct.
combined appears to be a combination of any of the above.
here's one answer from the web:
http://answers.yahoo.com/question/index?qid=20070726203353AAgDenC